In exercise 3, set up and evaluate a double integral that represents the area of the region.
Answer #1

The final answer will be 12\pi .

We will be making use of polar coordinates in order to evaluate the above integral. That is, rather than x and y, we will make use of r and \theta where,

x=r\cos{\theta}

and

y=r\sin{\theta}.

Range of r will be from 2 to 4, as clearly seen from the diagram. And range of \theta will be from 0 to 2\pi , since all the quadrants are colored.

We will set up the untequal uning cosuduates to make it sumplen to polar solve.. clucle r=2 and neue, auca Cuiside the lies o